OpenStudy (anonymous):

Give an example of a relation that is NOT a function and explain why it is not a function

OpenStudy (australopithecus):

so you are asking for an equation that is not a function?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

x^(2) + y^(2) = 1 is not a function because it does not pass the vertical line test, (use a search engine to figure out what the vertical line test is)

OpenStudy (australopithecus):

Here is the graph of the equation I just posted https://www.wolframalpha.com/input/?i=x^%282%29+%2B+y^%282%29+%3D+1 vertical line test basically states that if a vertical line can pass through the line of a graphed equation more than once anywhere on the graph the equation is not a function

OpenStudy (australopithecus):

the equation x^(2) + y^(2) = 1 would therefore not pass whereas the equation x + 1 = y would pass the vertical line test https://www.wolframalpha.com/input/?i=+x+%2B+1+%3D+y

OpenStudy (australopithecus):

notice you an draw a vertical line anywhere on that graph and it will only pass the equations line once thus it is a function

OpenStudy (australopithecus):

if you can acquire multiple y values for a single x in an equation then the equation is not a function, in the equation, x^(2) + y^(2) = 1 you can acquire 2 y values for every one x value therefore it is not a function

OpenStudy (australopithecus):

hope that helps